An airline finds that with probability 0.04 every passenger that makes a reservation on a particular flight will not show up. Consequently, their policy is to sell 200 reserved seats on a plane that has only 196 seats. (a) Find the probability that every person who shows up for the flight will find a seat available. (b) Use the Poisson approximation to compute this probability and compare your answer to the exact answer that you found in part (a).
Answer #1

a)

probability that every person who shows up for the flight will find a seat available

=P(at least 4 people not show up)=1-P(at most 3 people not show up)

=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)

=1-200C0(0.04)0(0.96)200 -200C1(0.04)1(0.96)199 -200C2(0.04)2(0.96)198 -200C3(0.04)3(0.96)197

=1-0.0395=0.9605

b)

from poisson distribution ; expected number of people not arrive =np=200*0.04=8

hence from Poisson approximation; probability that every person who shows up for the flight will find a seat available =1-P(X=0)-P(X=1)-P(X=2)-P(X=3)

=1-e-8*80/0!-e-8*81/1!-e-8*82/2!-e-8*83/3! =1-0.0424=0.9576